• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name averett

    Log entry time 20:06:00 on March 1,2006

    Entry number 164675

    keyword=Day Shift Report

    Day Shift Summary

    8:00 NMR measurement, 40%
    8:27-8:50 Beam position studies in empty position (0.6 uA, raster off). Results show no DC trips.
    8:50-9:00 Beams position studies in pol. 3He position (0.6 uA, raster on). DC trips in all measured positions except nominal "center" position.
    9:00 Beam off for beam studies. Access for laser work and to check flow on helium bags.
    13:00 Beam being restored.
    13:30-14:40 Beam position/rate studies using all targets with raster on.
    15:00-16:00 Beam quality measurements using HARP, Compton, Synchrotron measurements.